# Differentially Private Clustering in Data Streams

### v2

@@ 2 Preliminaries
Norms and heavy hitters. Let $p\geq 1$, the $\ell_{p}$-norm of a vector $\mathbf{x}=(x_{1},\ldots,x_{t})$ is defined as $\|\mathbf{x}\|_{p}=(\sum^{t}_{i=1}|x_{i}|^{p})^{1/p}$. Given a multiset ${\mathcal{S}}$, denote the frequency of an item $x$ appearing in ${\mathcal{S}}$ as $f(x)$. We say that an item $x$ is an $\theta$-heavy hitter ($\theta$-HH for short) if $f(x)\geq\theta\|{\mathcal{S}}\|_{1}$.
Differential Privacy. Streams ${\mathcal{S}}=(x_{1},\ldots,x_{T})$ and ${\mathcal{S}}^{\prime}=(x^{\prime}_{1},\ldots,x^{\prime}_{T})$ are neighboring if there exists at most one timestamp $t^{*}\in[T]$ for which $x_{t^{*}}\neq x^{\prime}_{t^{*}}$ and $x_{t}=x^{\prime}_{t}$ for all $t\neq t^{*}$.
@@
Let $\varepsilon>0$, $\gamma_{h}\in(0,0.5)$, $10<\theta<$ 1, $\xi\in(0,0.5)$. There is an $\varepsilon$-DP algorithm in the streaming continual release model such that with probability at least $1-\xi$, it always outputs a set $H\subseteq{\mathcal{U}}$ and a function $\hat{f}:H\to{\mathbb{R}}$ for every timestamp $t\in[T]$ such that
@@
Clustering. For points $x,y\in{\mathbb{R}}^{d}$, we let $d(x,y)=\|x-y\|_{2}$ be the Euclidean distance between $x$ and $y$. Given a set ${\mathcal{C}}$, we define $d(x,{\mathcal{C}}):=\min_{c\in{\mathcal{C}}}d(x,c)$.
For a set of centers $C$, we define the cost of clustering for the set ${\mathcal{S}}$ wrt $C$ as
$cost(C,{\mathcal{S}})=\sum_{x\in{\mathcal{S}}}d^{z}(x,C)$
where $z=1$ for $k$-median, and $z=2$ for $k$-means.
Our goal in DP clustering is to produce a set of $k$ centers $C_{\mathcal{S}}$ for input stream ${\mathcal{S}}$ such that (1) $C_{\mathcal{S}}$ is $(\varepsilon,\delta)$-DP wrt ${\mathcal{S}}$, and (2) $cost(C_{\mathcal{S}},{\mathcal{S}})\leq\alpha\cdot cost(C^{opt}_{\mathcal{S}},{\mathcal{S}})+\beta$.
Definition 2 ($(\kappa,\eta)$-coreset).
Given a point set $P$ in ${\mathbb{R}}^{d}$, a weighted subset $Q\subseteq P$ is a $(\kappa,\eta)$-coreset of $P$ for $k$-clustering ($k$-means or $k$-median) if
$\frac{1}{\kappa}\cdot\textsf{cost}(C,P)-\eta\leq\textsf{cost}(C,Q)\leq\kappa\cdot\textsf{cost}(C,P)+\eta$
for any set of $k$-centers $C\subseteq{\mathbb{R}}^{d}$.
In particular, given a point set $P$ in ${\mathbb{R}}^{d}$, a weighted subset $Q\subseteq P$ is a $(1+\gamma,\eta)$-coreset of $P$ for $k$-clustering ($k$-means or $k$-median) if
$(1-\gamma)\cdot\textsf{cost}(C,P)-\eta\leq\textsf{cost}(C,Q)\leq(1+\gamma)\cdot\textsf{cost}(C,P)+\eta$
for any set of $k$-centers $C\subseteq{\mathbb{R}}^{d}$.
Definition 3 ($(\kappa,\eta_{1},\eta_{2})$-semicoreset).
Given a point set $P$ in ${\mathbb{R}}^{d}$, a weighted subset $Q\subseteq P$ is a $(\kappa,\eta_{1},\eta_{2})$-semicoreset of $P$ for $k$-clustering ($k$-means or $k$-median) if
for any set of $k$-centers $C\subseteq{\mathbb{R}}^{d}$.
@@
Let ${\mathcal{A}}$ be an $(\varepsilon,\delta)$-DP (static) algorithm that takes as input a dataset $P$ of size at least $\Omega(k\cdot\varepsilon^{-1}\log(n))$ contained in a ball ${\mathcal{B}}(0,\Lambda)$ (with known center) and outputs a set of $k$ centers ${\mathcal{C}}$ contained in ${\mathcal{B}}$ such that $cost({\mathcal{C}},{\mathcal{X}})\leq O(1)cost({\mathcal{C}}^{opt},{\mathcal{X}})+V(n,d,k,\varepsilon,\delta)\cdot\Lambda^{p}$. Then for any fixed $\eta>0$ there exists a $(3\varepsilon,\delta)$-DP (sequential) algorithm ${\mathcal{B}}$ on the dataset ${\mathcal{Y}}$ that outputs a set $\hat{{\mathcal{P}}}\subset B$ such that for any set ${\mathcal{C}}$ of size at most $k$.
$cost(\hat{P},{\mathcal{C}})$ $\leq O(1)\cdot\left(cost({\mathcal{C}}^{opt}_{P},P)+cost({\mathcal{C}},P)+(V(n,d,k,\varepsilon,\delta)+k\varepsilon^{-1}\log(n))\cdot\Lambda^{p}\right)$ (3)
$cost(P,{\mathcal{C}})$ $\leq O(1)\cdot\left(cost({\mathcal{C}}^{opt}_{P},P)+cost({\mathcal{C}},\hat{P})+(V(n,d,k,\varepsilon,\delta)+k\varepsilon^{-1}\log(n))\cdot\Lambda^{p}\right)$ (4)
@@
Remark 2.
Any $(\kappa,\eta)$-coreset is also a $(\kappa,0,\eta)$-semicoreset. The output of any DP $k$-clustering algorithm can be made into a DP semicoreset via Theorem 7.

### v3

@@ Appendix A Preliminaries
Let $p\geq 1$, the $\ell_{p}$-norm of a vector $\mathbf{x}=(x_{1},\ldots,x_{t})$ is defined as $\|\mathbf{x}\|_{p}=(\sum^{t}_{i=1}|x_{i}|^{p})^{1/p}$. Given a multiset $\mathcal{S}$}, denote the frequency of an item $x$ appearing in $\mathcal{S}$} as $f(x)$. We say that an item $x$ is an $\theta$-heavy hitter ($\theta$-$H$ for short) if $f(x)\geq\theta\|{\mathcal{S}}\|_{1}$.
Differential Privacy background.
We state some DP theorems that we frequently utilize below. For a more detailed background on DP refer to [26].
@@
Let $\varepsilon>0$, $\gamma_{h}\in(0,0.5)$, $0<\theta<1$, $\xi\in(0,0.5)$. There is an $\varepsilon$-DP algorithm in the streaming continual release model such that with probability at least $1-\xi$, it always outputs a set $H\subseteq{\mathcal{U}}$ and a function $\hat{f}:H\to{\mathbb{R}}$ for every timestamp $t\in[T]$ such that
@@
Remark.
In algorithm 3 we use the notation Initialize and Update for the Binary Mechanism (BM) and DP-Heavy Hitters (DP-HH) primitives. For clarity:
  • Initialize sets up the respective data structure with its privacy budget and parameters (e.g., stream length, thresholds).
  • Update processes a single incoming item (or the empty symbol $\perp$) and returns the current private estimate. For BM this is a differentially private count, while for DP-HH it is a set of candidate heavy hitters together with approximate frequencies.
We do not re-write these algorithms explicitly, since our use is a direct adaptation of the continual release algorithms already stated in Theorem 9, Theorem 10 and in the cited references.
Useful Clustering Theorems.
We present some useful clustering theorems that we frequently utilize below.
We clarify that the statement below concerns sequential algorithms in the MPC model. However, when the entire input resides on a single machine, a sequential algorithm is trivially static. Since our use case involves static algorithms in such a setting, we restate the lemma accordingly to reflect this context more accurately.
@@
Let $\mathcal{A}$} be an $(\varepsilon,\delta)$-DP (static) algorithm that takes as input a dataset $P$ of size at least $\Omega(k\cdot\varepsilon^{-1}\log(n))$ contained in a ball ${\mathcal{B}}(0,\Lambda)$ (with known center) and outputs a set of $k$ centers $\mathcal{C}$} contained in $\mathcal{B}$} such that $cost({\mathcal{C}},{\mathcal{X}})\leq O(1)cost({\mathcal{C}}^{opt},{\mathcal{X}})+V(n,d,k,\varepsilon,\delta)\cdot\Lambda^{p}$. Then for any fixed $\eta>0$ there exists a $(3\varepsilon,\delta)$-DP (sequential) algorithm $\mathcal{B}$} on the dataset $\mathcal{Y}$} that outputs a set $\hat{P}\subset B$ such that for any set $\mathcal{C}$} of size at most $k$.
[…]
$cost({\mathcal{C}},P)$ $\leq O(1)\cdot\left(cost({\mathcal{C}}^{opt}_{P},P)+cost({\mathcal{C}},\hat{P})+(V(n,d,k,\varepsilon,\delta)+k\varepsilon^{-1}\log(n))\cdot\Lambda^{p}\right)$ (2)